OpenStudy (anonymous):

I need fast help please!! If vodka is 9.81 M in ethanol what is the molarity of ethanol in a drink that was made by mixing 47.0 mL of vodka in enough fruit juice to make 6.06 fl oz ( 1 fl oz =29.33 mL

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

What is the total volume of the drink?

OpenStudy (anonymous):

I don't know. It wasn't given.

OpenStudy (unklerhaukus):

it is given in fl oz, but what is it in mL?

OpenStudy (anonymous):

177.47 mL

OpenStudy (anonymous):

Sorry if I seem rushy. It's a timed thing

OpenStudy (unklerhaukus):

177.74 mL

OpenStudy (unklerhaukus):

how many mole is 47.0 mL of vodka (9.81 M)

OpenStudy (anonymous):

47.0 right?

OpenStudy (unklerhaukus):

47 moles?

OpenStudy (unklerhaukus):

way too much

OpenStudy (anonymous):

Isn't the 1 mol = 1 mL thing? Or am I thinking wrong?

OpenStudy (unklerhaukus):

concentration \(c\), is the number of moles \(n\), per volume \[c\,[\text M] = \frac{n[\text{mol}]}{V[\text L]}\]

OpenStudy (unklerhaukus):

if the concentration is 9.81 M it means there is 9.81 moles per litre

OpenStudy (anonymous):

Oh the molarity

OpenStudy (unklerhaukus):

the volume of (9.81M) vodka is 47mL the number of moles is n = 9.81 [mol/L] * 0.047 [L] =

OpenStudy (anonymous):

.46107

OpenStudy (unklerhaukus):

yes, 0.461 [mol]

OpenStudy (unklerhaukus):

now the volume becomes 177.74 mL, but the number of moles vodka stays the same

OpenStudy (unklerhaukus):

c = n/v = 0.461 [mol] / 177.74 mL = . . . [mol/mL] = .... [mol/L] = .... [M]

OpenStudy (anonymous):

.00259

OpenStudy (unklerhaukus):

always have units

OpenStudy (anonymous):

.00259 M

OpenStudy (unklerhaukus):

wrong units

OpenStudy (unklerhaukus):

0.00260 is [mol/mL] but you want [mol/L] = [M]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

How would I convert it to L?

OpenStudy (unklerhaukus):

1[L] = 1000 [mL] 1 = 1000 [mL/L] 0.00260 [mol/mL] * 1000 [mL/L] =

OpenStudy (anonymous):

2.6 M

OpenStudy (unklerhaukus):

right

OpenStudy (anonymous):

And that's all there is to it?

OpenStudy (unklerhaukus):

the easier way is 9.81 * 47/177

OpenStudy (anonymous):

That is easier

OpenStudy (anonymous):

Thanks

OpenStudy (unklerhaukus):

c_1 = n/V_1 c_2 = n/V_2 c_2 = c_1 * V_1/V_2

OpenStudy (anonymous):

So it' m1 v1 = m2 v2

OpenStudy (anonymous):

That sort of thing?

OpenStudy (unklerhaukus):

that is conservation of momentum, this is conservation of ethanol

OpenStudy (anonymous):

Okay.